Carla is 5 years old and Jim is 13 years younger than Peter. One year ago, Peter's age was twice the sum of Carla's & Jim's age. Find the present age of each one of them.

Answer by ankor@dixie-net.com(22740)

Carla is 5 years old and Jim is 13 years younger than Peter.
c = 5
j = p - 13
or
p = j + 13
:
One year ago, Peter's age was twice the sum of Carla's & Jim's age.
one year ago c = 4
p - 1 = 2[j-1+4]
p - 1 = 2(j+3)
p - 1 = 2j + 6
p = 2j + 6 + 1
p = 2j + 7
replace p with (j+13)
j + 13 = 2j + 7
13 - 7 = 2j - j
j = 6 yrs old is Jim
then
p = 6 + 13
p = 19 yrs old is Peter



Find the present age of each one of them. c=5, p=19, j=6